# Experimental Evidence of Néel-order-driven Magneto-optical Kerr Effect in an Altermagnetic Insulator

### v1

@@ Abstract
Introduction—Hematite ($\rm\alpha$-$\rm Fe_{2}O_{3}$) is a typical collinear antiferromagnetic (AFM) insulator at room-temperature, exhibiting weak ferromagnetism arising from the Dzyaloshinskii-Moriya (DM) interaction[1960_Moriya_PRL, 1960_Moriya_PR]. Hematite serves as a model system for exploring intriguing spintronic phenomena in antiferromagnetic insulators [2020_Cheng_PRL_FEO_switching, 2021_Cogulu_PRB_FEO_switching, 2022_Zhang_PRL_FEO_SOT, 2018_Lebrun_Nat_FEO_transport_domain, 2020_Han_FEO_magnon], owing to its distinctive properties such as high Néel temperature [1994_Morrish_Book_FEO], low damping constant [2018_Lebrun_Nat_FEO_transport_domain, 2020_Han_FEO_magnon], low magnetocrystalline anisotropy [2020_Cheng_PRL_FEO_switching] and large magneto-optical (MO) response [2021_Cogulu_PRB_FEO_switching]. In addition to the conventional second-order MO (Cotton-Mouton) effect typical in collinear AFMs[1969_Pisarev_MLB_1, 1969_Pisarev_MLB_2], hematite displays an unexpectedly large first-order MO signal that is comparable to that in ferromagnets, despite possessing a much weaker net magnetization[1971_Pisarev_MO_0001, 1973_Krinchik_MO_Longitudinal, 1974_Krinchik_MO_Sweep]. This phenomenon has been extensively investigated by spectroscopy [1973_Krinchik_MO_Longitudinal, 1981_Zubov_MO_spec, 2020_Ivantsov_MCD] and AFM domain imaging in hematite [1990_Appel_FEO_domain, 2018_Lebrun_Nat_FEO_transport_domain]. A widely cited but qualitative explanation attributes the anomalous MOKE in hematite to the Néel order under low-symmetry crystal environments, rather than to net magnetization or external magnetic fields—though direct evidence supporting this scenario has remained limited[1994_Morrish_Book_FEO, 2024_Kimel_JMMM_alterm_MO]. In early experiments aimed at verifying the Néel-order contribution in hematite, the measured MOKE signal contained entangled contributions from surface and volume, making it difficult to reveal the intrinsic response of hematite [1974_Krinchik_MO_Sweep]. More importantly, the contribution of the external magnetic field was completely overlooked without proper verification, resulting in experimental data inadequate to support the proposed explanation [1974_Krinchik_MO_Sweep, 1981_YFeO3]. In the theoretical aspect, it had been challenging to understand how a Néel order—with zero net magnetization—can give rise to a MOKE response. Consequently, a definitive conclusion regarding the origin of MOKE in hematite, as well as a microscopic understanding of the underlying mechanism, remains elusive.
@@
The recent discovery of altermagnetism offers a compelling new framework for understanding this anomalous magneto-optical behavior [2022_Smejkal_PRX_alterm]. Altermagnets, novel collinear AFMs characterized by spin splitting in $k$-space, have become a frontier topic due to their rich physical properties and great potential for spintronic applications [2022_Smejkal_PRX_alterm, 2025_Song_NatRev_altermag]. The intrinsic breaking of combined $\hat{P}\hat{T}$ symmetry in this material gives rise to ferromagnetic-like effects absent in conventional AFMs, including Néel-order-origin anomalous Hall effect (AHE) [2020_Smejkal_SciAdv_alterm_AHE, 2022_Feng_NatElec_RuO2_AHE, 2023_Gonzalez_PRL_MnTe_AHE, 2024_Takagi_NatMat_FeS_AHE], spin current generation [2021_Gonzalez_PRL_alterm, 2023_Bai_PRL_RuO2_spin] and X-ray magnetic circular dichroism (XMCD) [2023_Lovesey_PRB_MnTe_alterm, 2024_Hariki_PRL_MnTe_XMCD, 2024_Amin_Nat_MnTe_mapping]. Among these, the AHE is usually described by transverse conductivity ($\vec{\sigma}^{A}$), which is an axial vector representing the antisymmetric part of conductivity tensor and results in the generation of anomalous Hall signal. As the optical counterpart of the AHE, a MOKE response originating from the Néel order is similarly anticipated in altermagnets [2004_Yao, 2024_Kimel_JMMM_alterm_MO]. Notably, hematite is theoretically classified as a $g$-wave altermagnetic insulator [2022_Smejkal_PRX_alterm, 2024_Verbeek_PRR_FEO_alterm], which has been verified in experiments through XMCD and AHE measurements [2025_Galindez_AM]. Recent observation of magnon band splitting in hematite [2023_Kanj_SciAdv_magnon, 2025_Chen_PRL_FEO_magnon] may be also related to its altermagnetism. However, direct experimental evidence linking the MOKE in hematite—or in any altermagnetic insulator—to the Néel order remains absent.
@@
The primary challenge in elucidating the origin of the MOKE in hematite arises because the transverse optical conductivity $\vec{\sigma}^{A}$ could have contributions from three distinct sources [2024_Kimel_JMMM_alterm_MO, 2024_Xiao_altermag]: Néel vector $\vec{N}$, net magnetization $\vec{M}$, and external magnetic field $\vec{H}$, given
@@
where $\bm{\alpha}$, $\bm{\beta}$ and $\bm{\gamma}$ are rank-2 tensors, while the MOKE signal offers only two measurable quantities—the zero-field intercept $\sigma^{A}_{i,0}$ and the field slope $\partial\sigma^{A}_{i}/\partial H_{i}$—as illustrated in FIG. 1a, making it difficult to disentangle the underlying contributions [2024_Kimel_JMMM_alterm_MO]. In Eq. 1, only first order contributions are considered. Specifically, in collinear AFMs, $\vec{M}$ includes two distinct components: spontaneous magnetization ($M_{S}$) and field-induced magnetization ($M_{H}$). The MO contribution from $M_{S}$ remains constant in a single-domain state and is thus inseparable from the Néel-vector contribution within the intercept $\sigma^{A}_{i,0}$. Conversely, the MO contribution from $M_{H}$ is linearly proportional to $H$ provided the applied field remains significantly below the exchange field (approximately 970 T for hematite [1994_Morrish_Book_FEO]), thus blending with the $\vec{H}$ contribution in the slope $\partial\sigma^{A}_{i}/\partial H_{i}$. Consequently, with three unknowns in only two equations, we cannot confirm the existence of the Néel-order contribution. This issue is quite common in collinear antiferromagnets [1974_Krinchik_MO_Sweep, 1981_YFeO3]. To address this issue, previous AHE studies on FeS assumed negligible magnetization contribution based on ab-initio calculations, enabling clear separation of Néel and external-field contributions across temperature ranges[2024_Takagi_NatMat_FeS_AHE]. Therefore, the idea that the MOKE signal in hematite originates from the Néel order has remained a hypothesis. Whether this signal contains a significant contribution from net magnetization is also unresolved. To address this difficulty, we take advantage of the photon energy—a unique parameter in optical measurement—to resolve these problem. By systematically analyzing the photon-energy-dependent behavior of both intercept and slope across hematite samples with varied orientations, we rigorously identify the Néel order as the dominant contributor to the MOKE response, which is a hallmark feature of altermagnetism.
@@
Symmetry Analysis—Before presenting the results of the MOKE measurement, we perform a brief symmetry analysis on hematite with three typical Néel order orientations to clarify $\vec{\sigma}^{A}$ originated from altermagnetism. The Cartesian axes xyzxyz used here is defined with respect to the hexagonal lattice of hematite, as shown in FIG. 1b. Hematite has space group of $R\bar{3}c$ and spin arrangement of +++--+ in rhombohedral lattice [1994_Morrish_Book_FEO, 2008_Hill_FEO_ND]. For simplification, we consider spin orderings with zero-canting in hematite. For $\vec{N}\parallel x$ (FIG. 1c), the glide plane $g$ preserves the crystal structure but reverses $\vec{N}$. Consequently, the system exhibits symmetry under time reversal $\hat{T}$ followed by $g$, which constrains $\vec{\sigma}^{A}$ parallel to $g$, since $\vec{\sigma}^{A}\parallel g$ keeps invariant under the $g\hat{T}$ operation while $\vec{\sigma}^{A}\perp g$ reverses. For $\vec{N}\parallel y$ (FIG. 1d), $g$ preserves both the crystal structure and $\vec{N}$, therefore confining $\vec{\sigma}^{A}$ perpendicular to $g$. When $\vec{N}$ orients along the $c$-axis (FIG. 1e), the simultaneous preservation of $C_{3}$ and $C_{2}$ rotational symmetry forces $\vec{\sigma}^{A}$ to be parallel to both rotation axes and therefore vanishing. Taking all the cases into consideration, $\bm{\alpha}\cdot\vec{N}$ can be simplified as $\alpha\cdot(\vec{c}\times\vec{N})$ [2024_Xiao_altermag], where $\alpha$ is a scalar and $\vec{c}$ is the unit vector in $c$-axis. Although $\vec{\sigma}^{A}$ induced by $\vec{M}$ and $\vec{H}$ have no symmetry requirements, the spontaneous magnetization itself is relevant to crystal symmetry through DM interaction. As established[1994_Morrish_Book_FEO, 2024_Kimel_JMMM_alterm_MO, 2025_Roig_PRL], $\vec{M}_{S}$ can be expressed as $\vec{D}\times\vec{N}=D\cdot(\vec{c}\times\vec{N})$, where vector $\vec{D}$ parallel to the $c$-axis is the first order approximation of DM interaction. $\vec{M}_{H}$ can be expressed as $\mu\vec{H}$, since the magnetic susceptibilities in the $ab$-plane and along $c$-axis are nearly equal as reported [1994_Morrish_Book_FEO, 2023_Kanj_SciAdv_magnon]. Therefore, the total expression of transverse optical conductivity becomes:
@@
Sample Characterization—The experiments are conducted in hematite $(11\bar{2}0)$ and $(0001)$ bulk samples. X-ray diffraction characterizations confirm high monocrystallinity in both samples [supp]. The magnetic properties of samples in out-of-plane direction, corresponding to the $x$ and $z$ directions of hematite, are measured by vibrating sample magnetometer at room temperature, as shown in FIG. 2a and 2b. The $(11\bar{2}0)$ sample enters single domain state under magnetic field of 0.2 T along $x$-axis, indicating $ab$-plane as easy plane for $\vec{N}$ with low magnetocrystalline anisotropy. The $M_{S}$ of the sample in the $x$ direction is $2.13\times 10^{3}$ A/m. By adopting sublattice magnetization as $7.75\times 10^{5}$ A/m [2008_Hill_FEO_ND], the DM coefficient $D$ is determined as $1.37\times 10^{-3}$. The $(0001)$ sample shows near zero spontaneous magnetization in the $z$ direction. In this case, $\vec{N}$ lies along equivalent $1\bar{1}00]$ directions, which correspond to the relatively easy axes within the $ab$-plane. The magnetic susceptibilities are $1.22\times 10^{-3}$ along $x$-axis and $1.31\times 10^{-3}$ along $z$-axis, respectively. These results demonstrate good agreement with previous report [1994_Morrish_Book_FEO].
H-sweeping MOKE—The polar MOKE signal with out-of-plane magnetic field is measured for each sample, as illustrated in FIG. 2c and 2d. The procedure for the MOKE measurement follows previous works [1985_Querry_N, 1972_Johnson_N, 1998_You_formula, 2000_Qiu_MOreview], and the data analysis is included in the Supplemental Material [supp]. The incident angle of linearly polarized light to samples is set as 0.2 deg, which is a good approximation of normal incidence. The rotation and ellipticity of light are measured using the cross-polarizer method, and then converted to transverse optical conductivity $\sigma^{A}_{x}$ and $\sigma^{A}_{z}$ in $(11\bar{2}0)$ and $(0001)$ sample, respectively. According to Eq. 2, the dependence of $\sigma^{A}_{x}$ on $H_{x}$ can be expressed as follows:
@@
The field-sweeping results of $\sigma^{A}_{x}$ and $\sigma^{A}_{z}$ at the wavelength of 550 nm are shown in FIG. 2e and 2f, both of which exhibit near-zero slope above 0.2 T. The intercept of $\sigma^{A}_{x}$ in the single-domain region is $(-7.31+9.21i)\ \rm{S/cm}$ while that of $\sigma^{A}_{z}$ is $(0.06+0.08i)\ \rm{S/cm}$, corresponding to the description of Eq. 3 and Eq. 4. The large intercept corresponds to the magnitude of 1.7 mrad in MOKE signal, which is comparable to typical ferromagnets Fe and Co [2018_Higo_MOKE]. The slopes of $\sigma^{A}_{x}$ and $\sigma^{A}_{z}$ are $(4.15+5.88i)\times 10^{-2}\ \rm{S/(cm\cdot T)}$ and $(-4.34+6.18i)\times 10^{-2}\ \rm{S/(cm\cdot T)}$, which are below the noise level of the measurements.
@@
Theoretical Calculation—The dominance of the Néel-order contribution in the MOKE of hematite is further corroborated by our first-principles calculations on each contributions. The details of first-principles calculations and formula derivations are included in the Supplemental Material [supp] (see also references [1996_Ebert_RPP, 2004_Yao, 2016_Xiao_PRL, 2008_Mostofi_CPC, 1996_Kresse_PRB, 1999_Kresse_PRB, 2010_Book_Condensed, 2026_Xiao] therein). The anomalous mechanism of the MOKE is from the transverse optical conductivity. The general optical conductivity is defined as
@@
where $\sigma_{ij}(\omega)$ can be obtained using the Kubo-Greenwood formula [1996_Ebert_RPP, 2004_Yao, 2016_Xiao_PRL]. The transverse optical conductivity is then the antisymmetric part of $\sigma_{ij}(\omega)$, i.e., $\sigma^{A}_{i}=\frac{1}{2}\epsilon_{ijk}\sigma_{jk}$, where $\epsilon_{ijk}$ is the Levi-Civita symbol. Like the anomalous Hall effect, the transverse optical conductivity can be viewed as a spin-group symmetry breaking phenomenon [2025_Liu_PRX], and the Néel-order contribution occurs at linear order of the spin-orbit coupling (SOC) strength. In contrast, the $\vec{M}_{S}$ itself scales linearly with DM interaction, which is of the second order in the SOC [1960_Moriya_PRL, 1960_Moriya_PR], so its contribution to MOKE signal shall start at the third order. By constraining the direction of the local spin order in the first-principles calculations, we can calculate $\sigma_{i}^{A}(\omega)$ originating from the Néel order, $\alpha N$, and spontaneous magnetization, $\beta_{xx}M_{S}$, as shown in FIG. 3d and 3e. Clearly, the Néel-order contribution is much larger. The trend of $\alpha N$ matches well with the experimental values of $\sigma^{A}_{x,0}$. To further evaluate the magnetic-field induced MOKE, we consider the following second-order response function
@@
Summary—We clarify that the dominant contribution to the large MOKE signal in hematite is from the Néel order, which is a hallmark of the altermagnetic behavior in this material. The contributions from both the net magnetization and the external magnetic field are found to be negligible within the visible light range, which is verified in experiment and first-principles calculation for the first time. Our investigation demonstrates the ability of altermagnets to generate large MO signal, opening rich opportunities in the emergent field of altermagnetic opto-spintronics, including the exploration of altermagnetic insulator candidates like $\rm YFeO_{3}$, $\rm LaMnO_{3}$ and $\rm LaCrO_{3}$ [2018_Higo_MOKE, 2025_Naka_perovskites] and the development of altermagnet-based optical devices like zero-field optical isolators and integrated photonic logics gates [2022_rev_isolator, 2025_Lapointe_isolator, 2025_Pintus_logic].

### v2

@@ Abstract
Introduction—Hematite ($\rm\alpha$-$\rm Fe_{2}O_{3}$) is a typical collinear antiferromagnetic (AFM) insulator at room-temperature, exhibiting weak ferromagnetism arising from the Dzyaloshinskii-Moriya (DM) interaction[1, 2]. Hematite serves as a model system for exploring intriguing spintronic phenomena in antiferromagnetic insulators [3, 4, 5, 6, 7], owing to its distinctive properties such as high Néel temperature [8], low damping constant [6, 7], low magnetocrystalline anisotropy [3] and large magneto-optical (MO) response [4]. In addition to the conventional second-order MO (Cotton-Mouton) effect typical in collinear AFMs[9, 10], hematite displays an unexpectedly large first-order MO signal that is comparable to that in ferromagnets, despite possessing a much weaker net magnetization[11, 12, 13]. This phenomenon has been extensively investigated by spectroscopy [12, 14, 15] and AFM domain imaging in hematite [16, 6]. A widely cited but qualitative explanation attributes the anomalous MOKE in hematite to the Néel order under low-symmetry crystal environments, rather than to net magnetization or external magnetic fields—though direct evidence supporting this scenario has remained limited[8, 17]. In early experiments aimed at verifying the Néel-order contribution in hematite, the measured MOKE signal contained entangled contributions from surface and volume, making it difficult to reveal the intrinsic response of hematite [13]. More importantly, the contribution of the external magnetic field was completely overlooked without proper verification, resulting in experimental data inadequate to support the proposed explanation [13, 18]. In the theoretical aspect, it had been challenging to understand how a Néel order—with zero net magnetization—can give rise to a MOKE response. Consequently, a definitive conclusion regarding the origin of MOKE in hematite, as well as a microscopic understanding of the underlying mechanism, remains elusive.
@@
The recent discovery of altermagnetism offers a compelling new framework for understanding this anomalous magneto-optical behavior [19]. Altermagnets, novel collinear AFMs characterized by spin splitting in $k$-space, have become a frontier topic due to their rich physical properties and great potential for spintronic applications [19, 20]. The intrinsic breaking of combined $\hat{P}\hat{T}$ symmetry in this material gives rise to ferromagnetic-like effects absent in conventional AFMs, including Néel-order-origin anomalous Hall effect (AHE) [21, 22, 23, 24], spin current generation [25, 26] and X-ray magnetic circular dichroism (XMCD) [27, 28, 29]. Among these, the AHE is usually described by transverse conductivity ($\vec{\sigma}^{A}$), which is an axial vector representing the antisymmetric part of conductivity tensor and results in the generation of anomalous Hall signal. As the optical counterpart of the AHE, a MOKE response originating from the Néel order is similarly anticipated in altermagnets [30, 17]. Notably, hematite is theoretically classified as a $g$-wave altermagnetic insulator [19, 31], which has been verified in experiments through XMCD and AHE measurements [32]. Recent observation of magnon band splitting in hematite [33, 34] may be also related to its altermagnetism. However, direct experimental evidence linking the MOKE in hematite—or in any altermagnetic insulator—to the Néel order remains absent.
@@
The primary challenge in elucidating the origin of the MOKE in hematite arises because the transverse optical conductivity $\vec{\sigma}^{A}$ could have contributions from three distinct sources [17, 35]: Néel vector $\vec{N}$, net magnetization $\vec{M}$, and external magnetic field $\vec{H}$, given
@@
where $\bm{\alpha}$, $\bm{\beta}$ and $\bm{\gamma}$ are rank-2 tensors, while the MOKE signal offers only two measurable quantities—the zero-field intercept $\sigma^{A}_{i,0}$ and the field slope $\partial\sigma^{A}_{i}/\partial H_{i}$—as illustrated in FIG. 1a, making it difficult to disentangle the underlying contributions [17]. In Eq. 1, only first order contributions are considered. Specifically, in collinear AFMs, $\vec{M}$ includes two distinct components: spontaneous magnetization ($M_{S}$) and field-induced magnetization ($M_{H}$). The MO contribution from $M_{S}$ remains constant in a single-domain state and is thus inseparable from the Néel-vector contribution within the intercept $\sigma^{A}_{i,0}$. Conversely, the MO contribution from $M_{H}$ is linearly proportional to $H$ provided the applied field remains significantly below the exchange field (approximately 970 T for hematite [8]), thus blending with the $\vec{H}$ contribution in the slope $\partial\sigma^{A}_{i}/\partial H_{i}$. Consequently, with three unknowns in only two equations, we cannot confirm the existence of the Néel-order contribution. This issue is quite common in collinear antiferromagnets [13, 18]. To address this issue, previous AHE studies on FeS assumed negligible magnetization contribution based on ab-initio calculations, enabling clear separation of Néel and external-field contributions across temperature ranges[24]. Therefore, the idea that the MOKE signal in hematite originates from the Néel order has remained a hypothesis. Whether this signal contains a significant contribution from net magnetization is also unresolved. To address this difficulty, we take advantage of the photon energy—a unique parameter in optical measurement—to resolve these problem. By systematically analyzing the photon-energy-dependent behavior of both intercept and slope across hematite samples with varied orientations, we rigorously identify the Néel order as the dominant contributor to the MOKE response, which is a hallmark feature of altermagnetism.
@@
Symmetry Analysis—Before presenting the results of the MOKE measurement, we perform a brief symmetry analysis on hematite with three typical Néel order orientations to clarify $\vec{\sigma}^{A}$ originated from altermagnetism. The Cartesian axes xyzxyz used here is defined with respect to the hexagonal lattice of hematite, as shown in FIG. 1b. Hematite has space group of $R\bar{3}c$ and spin arrangement of +++--+ in rhombohedral lattice [8, 36]. For simplification, we consider spin orderings with zero-canting in hematite. For $\vec{N}\parallel x$ (FIG. 1c), the glide plane $g$ preserves the crystal structure but reverses $\vec{N}$. Consequently, the system exhibits symmetry under time reversal $\hat{T}$ followed by $g$, which constrains $\vec{\sigma}^{A}$ parallel to $g$, since $\vec{\sigma}^{A}\parallel g$ keeps invariant under the $g\hat{T}$ operation while $\vec{\sigma}^{A}\perp g$ reverses. For $\vec{N}\parallel y$ (FIG. 1d), $g$ preserves both the crystal structure and $\vec{N}$, therefore confining $\vec{\sigma}^{A}$ perpendicular to $g$. When $\vec{N}$ orients along the $c$-axis (FIG. 1e), the simultaneous preservation of $C_{3}$ and $C_{2}$ rotational symmetry forces $\vec{\sigma}^{A}$ to be parallel to both rotation axes and therefore vanishing. Taking all the cases into consideration, $\bm{\alpha}\cdot\vec{N}$ can be simplified as $\alpha\cdot(\vec{c}\times\vec{N})$ [35], where $\alpha$ is a scalar and $\vec{c}$ is the unit vector in $c$-axis. Although $\vec{\sigma}^{A}$ induced by $\vec{M}$ and $\vec{H}$ have no symmetry requirements, the spontaneous magnetization itself is relevant to crystal symmetry through DM interaction. As established[8, 17, 37], $\vec{M}_{S}$ can be expressed as $\vec{D}\times\vec{N}=D\cdot(\vec{c}\times\vec{N})$, where vector $\vec{D}$ parallel to the $c$-axis is the first order approximation of DM interaction. $\vec{M}_{H}$ can be expressed as $\mu\vec{H}$, since the magnetic susceptibilities in the $ab$-plane and along $c$-axis are nearly equal as reported [8, 33]. Therefore, the total expression of transverse optical conductivity becomes:
@@
Sample Characterization—The experiments are conducted in hematite $(11\bar{2}0)$ and $(0001)$ bulk samples. X-ray diffraction characterizations confirm high monocrystallinity in both samples [38]. The magnetic properties of samples in out-of-plane direction, corresponding to the $x$ and $z$ directions of hematite, are measured by vibrating sample magnetometer at room temperature, as shown in FIG. 2a and 2b. The $(11\bar{2}0)$ sample enters single domain state under magnetic field of 0.2 T along $x$-axis, indicating $ab$-plane as easy plane for $\vec{N}$ with low magnetocrystalline anisotropy. The $M_{S}$ of the sample in the $x$ direction is $2.13\times 10^{3}$ A/m. By adopting sublattice magnetization as $7.75\times 10^{5}$ A/m [36], the DM coefficient $D$ is determined as $1.37\times 10^{-3}$. The $(0001)$ sample shows near zero spontaneous magnetization in the $z$ direction. In this case, $\vec{N}$ lies along equivalent $1\bar{1}00]$ directions, which correspond to the relatively easy axes within the $ab$-plane. The magnetic susceptibilities are $1.22\times 10^{-3}$ along $x$-axis and $1.31\times 10^{-3}$ along $z$-axis, respectively. These results demonstrate good agreement with previous report [8].
H-sweeping MOKE—The polar MOKE signal with out-of-plane magnetic field is measured for each sample, as illustrated in FIG. 2c and 2d. The procedure for the MOKE measurement follows previous works [39, 40, 41, 42], and the data analysis is included in the Supplemental Material [38]. The incident angle of linearly polarized light to samples is set as 0.2 deg, which is a good approximation of normal incidence. The rotation and ellipticity of light are measured using the cross-polarizer method, and then converted to transverse optical conductivity $\sigma^{A}_{x}$ and $\sigma^{A}_{z}$ in $(11\bar{2}0)$ and $(0001)$ sample, respectively. According to Eq. 2, the dependence of $\sigma^{A}_{x}$ on $H_{x}$ can be expressed as follows:
@@
The field-sweeping results of $\sigma^{A}_{x}$ and $\sigma^{A}_{z}$ at the wavelength of 550 nm are shown in FIG. 2e and 2f, both of which exhibit near-zero slope above 0.2 T. The intercept of $\sigma^{A}_{x}$ in the single-domain region is $(-7.31+9.21i)\ \rm{S/cm}$ while that of $\sigma^{A}_{z}$ is $(0.06+0.08i)\ \rm{S/cm}$, corresponding to the description of Eq. 3 and Eq. 4. The large intercept corresponds to the magnitude of 1.7 mrad in MOKE signal, which is comparable to typical ferromagnets Fe and Co [43]. The slopes of $\sigma^{A}_{x}$ and $\sigma^{A}_{z}$ are $(4.15+5.88i)\times 10^{-2}\ \rm{S/(cm\cdot T)}$ and $(-4.34+6.18i)\times 10^{-2}\ \rm{S/(cm\cdot T)}$, which are below the noise level of the measurements.
@@
Theoretical Calculation—The dominance of the Néel-order contribution in the MOKE of hematite is further corroborated by our first-principles calculations on each contributions. The details of first-principles calculations and formula derivations are included in the Supplemental Material [38] (see also references [44, 30, 45, 46, 47, 48, 49, 50] therein). The anomalous mechanism of the MOKE is from the transverse optical conductivity. The general optical conductivity is defined as
@@
where $\sigma_{ij}(\omega)$ can be obtained using the Kubo-Greenwood formula [44, 30, 45]. The transverse optical conductivity is then the antisymmetric part of $\sigma_{ij}(\omega)$, i.e., $\sigma^{A}_{i}=\frac{1}{2}\epsilon_{ijk}\sigma_{jk}$, where $\epsilon_{ijk}$ is the Levi-Civita symbol. Like the anomalous Hall effect, the transverse optical conductivity can be viewed as a spin-group symmetry breaking phenomenon [51], and the Néel-order contribution occurs at linear order of the spin-orbit coupling (SOC) strength. In contrast, the $\vec{M}_{S}$ itself scales linearly with DM interaction, which is of the second order in the SOC [1, 2], so its contribution to MOKE signal shall start at the third order. By constraining the direction of the local spin order in the first-principles calculations, we can calculate $\sigma_{i}^{A}(\omega)$ originating from the Néel order, $\alpha N$, and spontaneous magnetization, $\beta_{xx}M_{S}$, as shown in FIG. 3d and 3e. Clearly, the Néel-order contribution is much larger. The trend of $\alpha N$ matches well with the experimental values of $\sigma^{A}_{x,0}$. To further evaluate the magnetic-field induced MOKE, we consider the following second-order response function
@@
Summary—We clarify that the dominant contribution to the large MOKE signal in hematite is from the Néel order, which is a hallmark of the altermagnetic behavior in this material. The contributions from both the net magnetization and the external magnetic field are found to be negligible within the visible light range, which is verified in experiment and first-principles calculation for the first time. Our investigation demonstrates the ability of altermagnets to generate large MO signal, opening rich opportunities in the emergent field of altermagnetic opto-spintronics, including the exploration of altermagnetic insulator candidates like $\rm YFeO_{3}$, $\rm LaMnO_{3}$ and $\rm LaCrO_{3}$ [43, 52] and the development of altermagnet-based optical devices like zero-field optical isolators and integrated photonic logics gates [53, 54, 55].
Note added: During the review process, we are aware that the work of J. Lou et al. [56] reaches similar conclusions.